OpenStudy (anonymous):

evaluate the trig integral: integral [0,pi/6] of sqrt(1+cos2x) dx

hartnn (hartnn):

you know the formulas for cos 2x ?

OpenStudy (anonymous):

I've found that it uses double angle formula - but I'm not familiar with it at all

OpenStudy (anonymous):

the fact that it is equal to three things confuses me and i don't know which to pick

hartnn (hartnn):

\(\Large \cos 2x = 2 \cos^2 x-1\) use this.

OpenStudy (anonymous):

Hmm...can you explain why you chose this one?

OpenStudy (anonymous):

Is there a rule of thumb I guess is what i'm asking, if it's all just something you learn from experience then i guess i'm s.o.l. for now

OpenStudy (anonymous):

to get rid of the 1

OpenStudy (anonymous):

You could have chosen any of them and it would have worked out.

hartnn (hartnn):

yes, i considered all 3 formulas, (remember we need the expression for "1+cos2x") cos2x = cos^2x - sin^2 x 1+ cos2x = 1+cos^2 x - sin^2x not much useful cos 2x = 1-2sin^2 x 1+cos2x = 2-2sin^2x hmm, can be used cos 2x = 2cos^2x -1 1+cos2x = 2cos^2 x very simple so that sqrt (1+cos 2x) is just sqrt 2 cos x

OpenStudy (anonymous):

If you use \(\cos(2x) = \cos^2(x)-\sin^2(x)\)Then you'd have: \[ 1+\cos^2(x)-\sin^2(x) = \cos^2(x)+1-\sin^2(x) =\cos^2(x)+\cos^2(x) = 2\cos^2(x) \]

OpenStudy (anonymous):

beautiful, thanks hartnn and wio :)

OpenStudy (anonymous):

I'll work on it from there